OpenStudy (anonymous):

[radical -2] -3 simplify using imaginary unit i

jimthompson5910 (jim_thompson5910):

By definition, \[\Large i = \sqrt{-1}\]

jimthompson5910 (jim_thompson5910):

We'll use the rule \[\Large \sqrt{x*y} = \sqrt{x}*\sqrt{y}\] to simplify

jimthompson5910 (jim_thompson5910):

use the rule and definition above to get... \[\Large \sqrt{-2} - 3\] \[\Large \sqrt{-1*2} - 3\] \[\Large \sqrt{-1}*\sqrt{2} - 3\] \[\Large i*\sqrt{2} - 3\] \[\Large -3 + (\sqrt{2})*i\] The last expression is in the form a+bi form where \(\Large a = -3\) and \(\Large b = \sqrt{2}\)